OpenStudy (anonymous):

Find the square. (7m - 3)^2

OpenStudy (anonymous):

(7m - 3)(7m - 3) and distribute. Or use the trinomial if you memorized it.

OpenStudy (anonymous):

\(49m^2 -42m +9\)

OpenStudy (anonymous):

never done this can you write it out to the answer

OpenStudy (anonymous):

It is pointless copying down answers, you will never learn this way.

OpenStudy (anonymous):

that's why i asked you to write it out? :)

OpenStudy (anonymous):

You should memorise these 3 formulas: \[(a+b)^2 = a^2 + 2ab + b^2\] \[(a-b)^2 = a^2 - 2ab + b^2\] \[(a+b)(a-b) = a^2 - b^2\]

OpenStudy (anonymous):

Apply formula 2 to this question. (7m - 3)^2 a = 7m b = -3 Is the asnwer obvious?

OpenStudy (anonymous):

Bd2zBone is right @WooTangClanANTFW

OpenStudy (anonymous):

so if the question is (2n + 2)(2n - 2) for finding the product do i use the third formula?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So 2n + 2 is 4n? right...

OpenStudy (anonymous):

or 4n^2 i dont see where the power comes from

OpenStudy (anonymous):

(2n + 2)(2n - 2) a = 2n b = 2 so what is a^2 and what is b^2?

OpenStudy (anonymous):

? so its 2n^2? and 2^2?

OpenStudy (anonymous):

\[(2n)^2 = 2^2n^2 = 4n^2\] \[2^2 = 4\] \[4n^2 - 4\]